a 2,4 high container in a chemical plant stores three separate chemicals ,one on top of each other .These chemicals have densities of 1450 kg/m^3,550kg/m^3 respectively. Each chemical has a height of 0,8 m. Determine the total force on the 1,5 m side of the container and where it acts


1
Expert's answer
2021-05-26T09:32:01-0400

"F=PA"

"A=S^2=1.5^2=2.25m^2"

"P=P_1+P_2+P_3"

"P=\rho_1h_1g+\rho_2h_2g+\rho_3h_3g=9.8(1450*2.4+550*1.6+200*0.8)=44296Nm"

"F=44296 *2.25=99666N"


The force acts on the bottom side of the container.